Find the orthogonal canonical reduction of the quadratic form
1
Expert's answer
2020-02-07T12:17:29-0500

Find the orthogonal canonical reduction of the quadratic form

"Q=x^2+y^2+z^2-2xy-2xz-2yz"


The quadratic form can be expressed in matrix notation as


"Q=\bold x^TA\bold x=\begin{pmatrix}
 x & y & z
\end{pmatrix}^T\begin{pmatrix}
 1 & -1 & -1 \\
 -1 & 1 & -1 \\
 -1 & -1 & 1
\end{pmatrix}\begin{pmatrix}
 x \\
 y \\
 z 
\end{pmatrix}"

The characteristic equation of the matrix "A" is


"\begin{vmatrix}
 1-\lambda & -1 & -1 \\
 -1 & 1-\lambda & -1 \\
 -1 & -1 & 1-\lambda
\end{vmatrix}=0"

"(1-\lambda)\begin{vmatrix}
 1-\lambda & -1 \\
 -1 & 1-\lambda
\end{vmatrix}-(-1)\begin{vmatrix}
 -1 & -1 \\
 -1 & 1-\lambda
\end{vmatrix}+""(-1)\begin{vmatrix}
 -1 & 1-\lambda \\
 -1& -1
\end{vmatrix}=0"


"(1-\lambda)((1-\lambda)^2-1)+(-1+\lambda-1)-(1+1-\lambda)=0"

"(1-\lambda)(1-2\lambda+\lambda^2-1)+\lambda-2-2+\lambda=0"

"\lambda(1-\lambda)(\lambda-2)+2(\lambda-2)=0"

"-(\lambda-2)(\lambda^2-\lambda-2)=0"

"(\lambda-2)(\lambda+1)(\lambda-2)=0"


"\lambda_1=\lambda_2=2, \lambda_3=-1" are eigenvalues.

"Q=2x'^2+2y'^2-z'^2" is the orthogonal canonical reduction.


"\lambda=2: \begin{pmatrix}
 1-\lambda & -1 & -1 \\
 -1 & 1-\lambda & -1 \\
 -1 & -1 & 1-\lambda
\end{pmatrix}=\begin{pmatrix}
 -1 & -1 & -1 \\
 -1 & -1 & -1 \\
 -1 & -1 & -1
\end{pmatrix}"

Perform row operations to obtain the rref of the matrix:

"R_2=R_2-R_1: \begin{pmatrix}
 -1 & -1 & -1 \\
 0 & 0 & 0 \\
 -1 & -1 & -1
\end{pmatrix}"

"R_3=R_3-R_1: \begin{pmatrix}
 -1 & -1 & -1 \\
 0 & 0 & 0 \\
 0 & 0 & 0
\end{pmatrix}"

"R_1=-1\cdot R_1: \begin{pmatrix}
 1 & 1 & 1 \\
 0 & 0 & 0 \\
 0 & 0 & 0
\end{pmatrix}"

Now, solve the matrix equation


"\begin{pmatrix}
 1 & 1 & 1 \\
 0 & 0 & 0 \\
 0 & 0 & 0
\end{pmatrix}
\begin{pmatrix}
 v_1 \\
 v_2 \\
 v_3
\end{pmatrix}=
\begin{pmatrix}
 0 \\
 0 \\
 0
\end{pmatrix}"

If we take "v_2=t, v_3=s," then "v_1=-t-s, v_2=t, v_3=s"

Therefore


"\bold v=\begin{pmatrix}
 -t-s \\
 t \\
 s
\end{pmatrix}=\begin{pmatrix}
 -1 \\
 1 \\
 0
\end{pmatrix}t+\begin{pmatrix}
 -1 \\
 0 \\
 1
\end{pmatrix}s"

"\lambda=-1: \begin{pmatrix}
 1-\lambda & -1 & -1 \\
 -1 & 1-\lambda & -1 \\
 -1 & -1 & 1-\lambda
\end{pmatrix}=\begin{pmatrix}
 2 & -1 & -1 \\
 -1 & 2 & -1 \\
 -1 & -1 & 2
\end{pmatrix}"

Perform row operations to obtain the rref of the matrix:

"R_1=R_1/2: \begin{pmatrix}
 1 & -1/2 & -1/2 \\
 -1 & 2 & -1 \\
 -1 & -1 & 2
\end{pmatrix}"

"R_2=R_2+R_1: \begin{pmatrix}
 1 & -1/2 & -1/2 \\
 0 & 3/2 & -3/2 \\
 -1 & -1 & 2
\end{pmatrix}"

"R_3=R_3+R_1: \begin{pmatrix}
 1 & -1/2 & -1/2 \\
 0 & 3/2 & -3/2 \\
 0 & -3/2 & 3/2
\end{pmatrix}"

"R_3=R_3+R_2: \begin{pmatrix}
 1 & -1/2 & -1/2 \\
 0 & 3/2 & -3/2 \\
 0 & 0 & 0
\end{pmatrix}"

"R_2=(2/3)R_2: \begin{pmatrix}
 1 & -1/2 & -1/2 \\
 0 & 1 & -1 \\
 0 & 0 & 0
\end{pmatrix}"

"R_1=R_1+(1/2)R_2: \begin{pmatrix}
 1 & 0 & -1 \\
 0 & 1 & -1 \\
 0 & 0 & 0
\end{pmatrix}"

Now, solve the matrix equation


"\begin{pmatrix}
 1 & 0 & -1 \\
 0 & 1 & -1 \\
 0 & 0 & 0
\end{pmatrix}
\begin{pmatrix}
 v_1 \\
 v_2 \\
 v_3
\end{pmatrix}=
\begin{pmatrix}
 0 \\
 0 \\
 0
\end{pmatrix}"

If we take "v_3=0," then "v_1=t,v_2=t,v_3=t."


"\bold v=\begin{pmatrix}
 t \\
 t \\
 t
\end{pmatrix}=\begin{pmatrix}
 1 \\
 1 \\
 1
\end{pmatrix}t"

"\sqrt {(-1)^2+1^2+0^2}=\sqrt{2}"

"\sqrt {(-1)^2+0^2+1^2}=\sqrt{2}"

"\sqrt {1^2+1^2+1^2}=\sqrt{3}"

"p_1={1 \over \sqrt{2}}\begin{pmatrix}
 -1 \\
 1 \\
 0
\end{pmatrix},p_2={1 \over \sqrt{2}}\begin{pmatrix}
 -1 \\
 0 \\
 1
\end{pmatrix},p_3={1 \over \sqrt{3}}\begin{pmatrix}
 1 \\
 1 \\
 1
\end{pmatrix}" are the principal axes.

"P=\begin{pmatrix}
 -{1 \over \sqrt{2}} & -{1 \over \sqrt{2}} & {1 \over \sqrt{3}} \\
 {1 \over \sqrt{2}} & 0 & {1 \over \sqrt{3}} \\
 0 & {1 \over \sqrt{2}} & {1 \over \sqrt{3}}
\end{pmatrix}"

Thus, a substitution "\bold x=P\bold x'" that eliminates the cross product terms is


"\begin{pmatrix}
 x \\
 y \\
 z 
\end{pmatrix}=\begin{pmatrix}
 -{1 \over \sqrt{2}} & -{1 \over \sqrt{2}} & {1 \over \sqrt{3}} \\
 {1 \over \sqrt{2}} & 0 & {1 \over \sqrt{3}} \\
 0 & {1 \over \sqrt{2}} & {1 \over \sqrt{3}}
\end{pmatrix}\begin{pmatrix}
 x' \\
 y' \\
 z' 
\end{pmatrix}"

This produces the new quadratic form

"Q=\bold x'^T(P^TAP)\bold x'="

"=\begin{pmatrix}
 x' & y' & z'
\end{pmatrix}^T\begin{pmatrix}
 2 & 0 & 0 \\
 0 & 2 & 0 \\
 0 & 0 & -1
\end{pmatrix}\begin{pmatrix}
 x' \\
 y' \\
 z' 
\end{pmatrix}="

"=2x'^2+2y'^2-z'^2"

in which there are no cross product terms.


"Q=2x'^2+2y'^2-z'^2" is the orthogonal canonical reduction.